OpenStudy (anonymous):

solve this by completing the square.. x2 + 4x – 12 = 0

hero (hero):

I don't know why no one helps you

OpenStudy (anonymous):

wolframalpha does http://www.wolframalpha.com/input/?i=x%5E2+%2B+4x+%E2%80%93+12+%3D+0+

OpenStudy (anonymous):

take half of 4, then square. (4/2)^2 = 4. That's the number that needs to be on the end x^2 + 4x +4 -4-12= 0 Then factor (x+2)^2 - 16 = 0 (x+2)^2 = 16 x+2 = +-4 x=-2+-(-4) = 2,-6

OpenStudy (saifoo.khan):

(x^2+4x) - 12 = 0 x^2 + 4x + (4/2)^2 -(4/2)^2 - 12 = 0 x^2 + 4x + 2^2 - (2)^2 - 12 = 0 (x+2)^2 - 4 - 12 = 0 (x+2)^2 - 16 = 0

hero (hero):